# Heterogeneous Multi-Agent Reinforcement Learning with Attention for Cooperative and Scalable Feature Transformation

### v1

@@ 1. Introduction
Feature transformation creates more informative features by mathematically transforming the original ones. This enhances pattern distinguishability and improves model performance. Despite the remarkable success of deep learning, feature transformation remains essential for structured data (e.g., tabular data) (fu2024tabular) and scenarios requiring high interpretability or the integration of domain-specific knowledge. For instance, in finance, features like debt-to-income ratio and credit utilization are derived from raw data. These features enhance interpretability and improve the performance of tasks such as credit risk assessment and loan approval. In healthcare, risk scores are derived by transforming clinical measurements into interpretable features. These features support more accurate diagnosis and treatment planning. All of these examples demonstrate that feature transformation still plays an important role in real-world scenarios and AI applications.
Existing work on automated feature transformation can be categorized: 1) Expansion reduction methods (kanter2015deep; horn2020autofeat; khurana2016cognito), which expand the feature space by applying mathematical operations and then select the most suitable features to reduce the feature space size; 2) Iterative feedback loops (wang2022group; khurana2018feature; tran2016genetic; huang2025collaborative; hu2024reinforcement), which iteratively refine features based on feedback from downstream tasks, optimized by RL or evolutionary algorithms; 3) AutoML-based methods (chen2019neural; zhu2022difer; gao2025gpt; liu2025continuous; wang2023reinforcement; xiao2023beyond), which frame feature transformation as a search problem for neural architectures and solve it using AutoML techniques. Among these, iterative-feedback paradigms excel at dynamically refining features over iterations. They can adapt to diverse downstream tasks and often yield effective feature spaces.
Recent studies have shown the effectiveness of RL in feature transformation (wang2022group; hu2024reinforcement; huang2024enhancing; ying2023self; xiao2024traceable; ying2025survey; xiao2023traceable). Typically, RL agents are employed to select the necessary features and mathematical operations to generate new informative features. This process is repeated iteratively with the goal of optimizing downstream task performance, continuing until the maximum iteration limit is reached. However, existing RL-based methods still face two key challenges:
@@
Figure 1. Framework Overview. HAFT consists of three heterogeneous agents: two feature agents that select candidate features and one operation agent that determines a mathematical operation. A central critic leverages global information from the current feature space to coordinate the agents and enhance policy learning. This feature transformation process continues iteratively until an optimal feature set is identified or the iteration limit is reached.:
@@ 3.2.1. Multi-Head Attention-Based Feature Agent
Why dynamic evolving feature space matters? During the iterative process, the size of the feature space changes after each iteration. This dynamic nature makes it challenging to design an agent structure that can handle variable input sizes and capture complex feature interactions effectively. Prior research (wang2022group) employs group-wise clustering to address the dynamic changes within the feature space. While effective in capturing high-level group structures, this approach struggles with scalability and modeling fine-grained feature relationships. Thus, we propose a multi-head attention-based (vaswani2017attention) agent structure to address these limitations. In the following, we show the calculation process using the feature space at iteration $t$. To simplify notation, we omit $t$, using $\mathcal{F}$ to represent $\mathcal{F}_{t}$.
@@ 3.3. Enhancing Collaboration with a Shared Critic and Advantage Decomposition
Why a shared central critic matters? Existing works (wang2022group) on automated feature transformation rely on partial information (e.g., decisions from previous agents) to coordinate and learn transformation policies. However, convergence in this setting is particularly challenging, as each agent’s policy is influenced not only by preceding agents but also by subsequent ones, creating complex interdependencies. To improve communication among agents and learn better feature transformation policies, a shared central critic module is essential. This necessitates robust credit assignment and alignment mechanisms. Unlike conventional multi-agent systems or feature-selection RL—where action spaces remain fixed and counterfactual baselines are straightforward to define—feature transformation presents unique difficulties. The dynamic expansion of the feature pool throughout the learning process makes credit assignment substantially more complex, as each agent’s contribution becomes increasingly entangled with the evolving feature space. This contrasts with typical multi-agent settings where individual contributions to shared rewards are more readily identifiable or where counterfactual reasoning can be more directly applied.
To address these unique challenges, we apply a Shared Critic with multi-agent advantage decomposition and a sequential policy-update scheme(kuba2021trust; yu2022surprising). The critic consumes aggregated global information and outputs a single scalar value estimate:
@@
Figure 2. Two‑branch state encoder. A statistics branch expands the dynamic feature set and normalizes it to $\mathbb{R}^{1\times 49}$; an attention branch embeds features, applies multi‑head attention, and pooling to obtain $z_{attn}$. Finally create input $s^{\text{critic}}=[S_{\text{stats}},z_{\text{attn}}]$.:
@@ 3.4. State Encoding for Shared Critic
Figure 3. Influence of Shared Critic and Advantage Decomposition:
@@ 4.1.1. Datasets
We use 23 publicly available datasets from UCI (uci_dataset), LibSVM(libsvm_dataset), Kaggle(kaggle_dataset), and OpenML(openml_dataset) to conduct our experiments. The 23 datasets involve 14 classification tasks and 9 regression tasks. Table 1 reflects the statistics of the data.
@@ 4.1.2. Evaluation Metrics
We adopt standard evaluation metrics to ensure a comprehensive analysis of our model’s performance. Regression tasks were evaluated using the following metrics: 1-Relative Absolute Error (1-RAE) (wang2022group), 1-Mean Absolute Error (1-MAE), 1-Mean Squared Error (1-MSE) and coefficient of determination ($R^{2}$). For classification tasks, assessment was conducted using Precision, Recall, F1-score and Accuracy. The formulae for the F1-score and 1-RAE are given by: $F_{1}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ and $\text{1-RAE}=1-\frac{\sum_{i=1}^{n}|y_{i}-\tilde{y}_{i}|}{\sum_{i=1}^{n}|y_{i}-\bar{y}_{i}|},$ where $y_{i}$, $\tilde{y}_{i}$, and $\bar{y}_{i}$ represent the ground truth, predictions, and the mean of the ground truth respectively.
@@ 4.1.3. Baseline Model
We compare our method with 8 widely-used feature transformation methods. 1) RDG generates feature-operation-feature transformation records randomly to create a new feature space. 2) ERG applies operations to each feature to expand the feature space and selects the crucial features as the new feature set. 3) LDA (blei2003latent) is a matrix factorization-based method that obtains the factorized hidden state as the generated feature space. 4) AFAT (horn2020autofeat) is an enhanced version of ERG that repeatedly generates new features and uses multi-step feature selection to identify informative ones. 5) NFS (chen2019neural) models the transformation sequence of each feature and uses reinforcement learning (RL) to optimize the entire feature generation process. 6) TTG (khurana2018feature) formulates the transformation process as a graph and implements an RL-based search method to identify the best feature set. 7) GRFG (wang2022group) utilizes three collaborative reinforcement agents for feature generation and introduces a feature grouping strategy to accelerate agent learning. 8) DIFER (zhu2022difer) embeds randomly generated feature transformation records using a seq2seq model and employs gradient search to identify the best feature set. Besides that, to validate the necessity of each component of HAFT, we develop three model variants: a. HAFT-c replaces the shared critic in HAFT with separate critics for each agent; b. HAFT-a removes the advantage decomposition component from HAFT while retaining the shared critic. c. HAFT-u only utilize the $S_{stats}$ as the shared critic input.
@@ 4.1.4. Experimental Settings
Figure 4. Influence of Different State Encoding:
@@
Figure 5. Time Complexity Comparison between HAFT and GRFG:
@@ 4.1.5. Hyperparameters and Reproducibility
The operation set includes square root, square, cosine, sine, tangent, exponential, cube, logarithm, reciprocal, quantile transformation, min–max scaling, sigmoid, addition, subtraction, multiplication, and division. We cap the process at 100 episodes, each with up to 25 steps, and select the top 20 features using a mutual-information-based metric to evaluate downstream performance. The entropy coefficient ranges from 0.01 to 0.3, and we explore learning rates of 1e–3, 1e–4, and 1e–5 for both the feature agent and operation agent, the dimension for $z_{\text{attn}}$¡=49 depending on the dataset. In terms of network architecture, the feature agent employs two Transformer encoder layers and one MLP projection layer. Meanwhile, the shared critic consists of four MLP layers, while the operation agent comprises two MLP layers, each employing LeakyReLU as the activation function. The parameters of all baseline models are configured according to the default settings specified in the corresponding papers, ensuring consistency in evaluation settings (ying2024unsupervised).
@@ 4.2.2. Ablation Study 1: Shared Critic
Figure 6. Robustness Check of HAFT when confronted with different downstream ML models in terms of F1 score and 1-RAE :
@@
Figure 7. Case Study of Messidor_features dataset:
@@ 4.2.4. Robustness Check
This experiment aims to answer: Does HAFT exhibit robustness when confronted with various machine learning models serving as downstream tasks? To answer this question, we replace the downstream ML model with Random Forest (RF) (breiman2001random), XGBoost (XGB) (chen2016xgboost), Support Vector Machine (SVM) (cortes1995support), K-Nearest Neighbors (KNN) (cover1967nearest), and Ridge Regression (Ridge) (hoerl1970ridge), respectively. Figure 6 presents the comparison results in terms of F1 score or 1-RAE on Openml_586, Openml_618, Ionosphere, and SpectF datasets. We find that regardless of changes in downstream models, HAFT consistently refines the feature space and achieves better results than the original feature space. The underlying reason is that HAFT can customize feature transformation policies to the specific characteristics of downstream models, resulting in more effective feature spaces. Thus, this experiment demonstrates the robustness of HAFT across different downstream models.
@@ 4.2.8. Study of the Hyperparameter Sensitivity of HAFT
Figure 8. The hyperparameter sensitivity test on SpectF:
@@ 5. Related Works
Automated Feature Transformation aims to refine or augment the original feature space so that machine learning models can more effectively capture complex, high-order relationships among variables. Traditionally, feature transformation techniques fall into two broad categories: latent representation learning-based and feature transformation-based approaches. Latent representation learning-based methods, as demonstrated in  (bengio2013representation; guo2017deepfm), focus on learning an optimized, often high-dimensional latent space where feature interactions are implicitly encoded. While these methods are known to boost predictive accuracy, the resulting representations can be challenging to interpret or directly trace back to the original input space. In contrast, feature transformation-based methods (wang2022group; huang2024enhancing; hu2024reinforcement; xiao2024traceable; azim2024feature) generate new features using arithmetic or aggregate operations. Although these methods yield traceable features, they often encounter an exponentially growing search space problem and become inefficient when dealing with large feature sets. In contrast to previous approaches, our attention-based feature crossing strategy not only captures the inherent relationships among features but also scales efficiently. This is achieved by employing cascading agents guided by a shared critic to learn effective feature interaction policies.
@@
Reinforcement Learning investigates how intelligent agents should act within an environment to maximize the expected cumulative reward (sutton2018reinforcement; wiering2012reinforcement). Based on the learned policy, reinforcement learning algorithms can be classified into two main categories: value-based and policy-based. Value-based algorithms (mnih2013playing; van2016deep) estimate the value of states or state-action pairs to facilitate action selection, whereas policy-based algorithms (sutton1999policy) learn a probability distribution that maps states to actions. Furthermore, the actor-critic framework (schulman2017proximal) combines the strengths of both approaches. While single-agent reinforcement learning has seen remarkable success, recent work increasingly focuses on multi-agent settings (wen2022multi) where multiple agents collaborate toward a shared goal. However, variations in observations, actions, or rewards often arise, making effective communication critical for coordination and overall performance. Foerster et al. (foerster2016learning) introduced RIAL and DIAL to facilitate communication among agents. Building on this, Foerster et al. (foerster2018counterfactual) proposed the COMA, employing a centralized critic and a counterfactual baseline for credit assignment. This design supports effective coordination even when agents have distinct reward functions. Lowe et al. (lowe2017multi) extended the actor-critic paradigm by implementing a centralized critic for each agent and allowing them to learn continuous policies, improving the scalability and flexibility of MARL systems. Moreover, Lyu et al. (lyu2021contrasting) found that centralized critics do not consistently outperform their decentralized counterparts across various evaluation domains. Recent advances in the MARL field, such as the theoretical formulation and empirical adaptation of PPO to cooperative multi-agent settings (yu2022surprising; kuba2021trust), and Asynchronous Action Coordination (zhang2024sequential), have accelerated research progress in this area. Overall, these studies underscore the importance of communication, collaboration, and effective credit assignment in the context of MARL. In our work, we adopt a similar communication mechanism by introducing a customized cascading design alongside a shared critic to address our feature transformation task.

### v2

# arXiv:2511.21934v2[cs.LG] 27 Feb 2026

@@ 1. Introduction
Feature transformation creates more informative features by mathematically transforming the original ones. This enhances pattern distinguishability and improves model performance. Despite the remarkable success of deep learning, feature transformation remains essential for structured data (e.g., tabular data) (Fu et al., 2024) and scenarios requiring high interpretability or the integration of domain-specific knowledge. For instance, in finance, features like debt-to-income ratio and credit utilization are derived from raw data. These features enhance interpretability and improve the performance of tasks such as credit risk assessment and loan approval. In healthcare, risk scores are derived by transforming clinical measurements into interpretable features. These features support more accurate diagnosis and treatment planning. All of these examples demonstrate that feature transformation still plays an important role in real-world scenarios and AI applications.
Existing work on automated feature transformation can be categorized: 1) Expansion reduction methods (Kanter and Veeramachaneni, 2015; Horn et al., 2020; Khurana et al., 2016), which expand the feature space by applying mathematical operations and then select the most suitable features to reduce the feature space size; 2) Iterative feedback loops (Wang et al., 2022; Khurana et al., 2018; Tran et al., 2016; Huang et al., 2025; Hu et al., 2024), which iteratively refine features based on feedback from downstream tasks, optimized by RL or evolutionary algorithms; 3) AutoML-based methods (Chen et al., 2019; Zhu et al., 2022; Gao et al., 2025; Liu et al., 2025; Wang et al., 2023; Xiao et al., 2023b), which frame feature transformation as a search problem for neural architectures and solve it using AutoML techniques. Among these, iterative-feedback paradigms excel at dynamically refining features over iterations. They can adapt to diverse downstream tasks and often yield effective feature spaces.
Recent studies have shown the effectiveness of RL in feature transformation (Wang et al., 2022; Hu et al., 2024; Huang et al., 2024; Ying et al., 2023; Xiao et al., 2024; Ying et al., 2025; Xiao et al., 2023a). Typically, RL agents are employed to select the necessary features and mathematical operations to generate new informative features. This process is repeated iteratively with the goal of optimizing downstream task performance, continuing until the maximum iteration limit is reached. However, existing RL-based methods still face two key challenges:
@@
Figure 1. Framework Overview. HAFT consists of three heterogeneous agents: two feature agents that select candidate features and one operation agent that determines a mathematical operation. A central critic leverages global information from the current feature space to coordinate the agents and enhance policy learning. This feature transformation process continues iteratively until an optimal feature set is identified or the iteration limit is reached.
An overview of our framework
@@ 3.2.1. Multi-Head Attention-Based Feature Agent
Why dynamic evolving feature space matters? During the iterative process, the size of the feature space changes after each iteration. This dynamic nature makes it challenging to design an agent structure that can handle variable input sizes and capture complex feature interactions effectively. Prior research (Wang et al., 2022) employs group-wise clustering to address the dynamic changes within the feature space. While effective in capturing high-level group structures, this approach struggles with scalability and modeling fine-grained feature relationships. Thus, we propose a multi-head attention-based (Vaswani et al., 2017) agent structure to address these limitations. In the following, we show the calculation process using the feature space at iteration $t$. To simplify notation, we omit $t$, using $\mathcal{F}$ to represent $\mathcal{F}_{t}$.
@@ 3.3. Enhancing Collaboration with a Shared Critic and Advantage Decomposition
Why a shared central critic matters? Existing works (Wang et al., 2022) on automated feature transformation rely on partial information (e.g., decisions from previous agents) to coordinate and learn transformation policies. However, convergence in this setting is particularly challenging, as each agent’s policy is influenced not only by preceding agents but also by subsequent ones, creating complex interdependencies. To improve communication among agents and learn better feature transformation policies, a shared central critic module is essential. This necessitates robust credit assignment and alignment mechanisms. Unlike conventional multi-agent systems or feature-selection RL—where action spaces remain fixed and counterfactual baselines are straightforward to define—feature transformation presents unique difficulties. The dynamic expansion of the feature pool throughout the learning process makes credit assignment substantially more complex, as each agent’s contribution becomes increasingly entangled with the evolving feature space. This contrasts with typical multi-agent settings where individual contributions to shared rewards are more readily identifiable or where counterfactual reasoning can be more directly applied.
To address these unique challenges, we apply a Shared Critic with multi-agent advantage decomposition and a sequential policy-update scheme(Kuba et al., 2021; Yu et al., 2022). The critic consumes aggregated global information and outputs a single scalar value estimate:
@@
Figure 2. Two‑branch state encoder. A statistics branch expands the dynamic feature set and normalizes it to $\mathbb{R}^{1\times 49}$; an attention branch embeds features, applies multi‑head attention, and pooling to obtain $z_{attn}$. Finally create input $s^{\text{critic}}=[S_{\text{stats}},z_{\text{attn}}]$.
Central Critic Input State.
@@ 3.4. State Encoding for Shared Critic
Figure 3. Influence of Shared Critic and Advantage Decomposition
Effect of Shared Critic and Advantage Decomposition
@@ 4.1.1. Datasets
We use 23 publicly available datasets from UCI (UCI Machine Learning Repository, 2022), LibSVM(Lin, 2022), Kaggle(Howard, 2022), and OpenML(Public, 2022) to conduct our experiments. The 23 datasets involve 14 classification tasks and 9 regression tasks. Table 1 reflects the statistics of the data.
@@ 4.1.2. Evaluation Metrics
We adopt standard evaluation metrics to ensure a comprehensive analysis of our model’s performance. Regression tasks were evaluated using the following metrics: 1-Relative Absolute Error (1-RAE) (Wang et al., 2022), 1-Mean Absolute Error (1-MAE), 1-Mean Squared Error (1-MSE) and coefficient of determination ($R^{2}$). For classification tasks, assessment was conducted using Precision, Recall, F1-score and Accuracy. The formulae for the F1-score and 1-RAE are given by: $F_{1}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ and $\text{1-RAE}=1-\frac{\sum_{i=1}^{n}|y_{i}-\tilde{y}_{i}|}{\sum_{i=1}^{n}|y_{i}-\bar{y}_{i}|},$ where $y_{i}$, $\tilde{y}_{i}$, and $\bar{y}_{i}$ represent the ground truth, predictions, and the mean of the ground truth respectively.
@@ 4.1.3. Baseline Model
We compare our method with 8 widely-used feature transformation methods. 1) RDG generates feature-operation-feature transformation records randomly to create a new feature space. 2) ERG applies operations to each feature to expand the feature space and selects the crucial features as the new feature set. 3) LDA (Blei et al., 2003) is a matrix factorization-based method that obtains the factorized hidden state as the generated feature space. 4) AFAT (Horn et al., 2020) is an enhanced version of ERG that repeatedly generates new features and uses multi-step feature selection to identify informative ones. 5) NFS (Chen et al., 2019) models the transformation sequence of each feature and uses reinforcement learning (RL) to optimize the entire feature generation process. 6) TTG (Khurana et al., 2018) formulates the transformation process as a graph and implements an RL-based search method to identify the best feature set. 7) GRFG (Wang et al., 2022) utilizes three collaborative reinforcement agents for feature generation and introduces a feature grouping strategy to accelerate agent learning. 8) DIFER (Zhu et al., 2022) embeds randomly generated feature transformation records using a seq2seq model and employs gradient search to identify the best feature set. Besides that, to validate the necessity of each component of HAFT, we develop three model variants: a. HAFT-c replaces the shared critic in HAFT with separate critics for each agent; b. HAFT-a removes the advantage decomposition component from HAFT while retaining the shared critic. c. HAFT-u only utilize the $S_{stats}$ as the shared critic input.
@@ 4.1.4. Experimental Settings
Figure 4. Influence of Different State Encoding
Comparison of value function training curves under different state encodings on four datasets: Openml_589, Openml_620, German credit, and SVMGuide3.
@@
Figure 5. Time Complexity Comparison between HAFT and GRFG
Time Complexity between HAFT and GRFG
@@ 4.1.5. Hyperparameters and Reproducibility
The operation set includes square root, square, cosine, sine, tangent, exponential, cube, logarithm, reciprocal, quantile transformation, min–max scaling, sigmoid, addition, subtraction, multiplication, and division. We cap the process at 100 episodes, each with up to 25 steps, and select the top 20 features using a mutual-information-based metric to evaluate downstream performance. The entropy coefficient ranges from 0.01 to 0.3, and we explore learning rates of 1e–3, 1e–4, and 1e–5 for both the feature agent and operation agent, the dimension for $z_{\text{attn}}$¡=49 depending on the dataset. In terms of network architecture, the feature agent employs two Transformer encoder layers and one MLP projection layer. Meanwhile, the shared critic consists of four MLP layers, while the operation agent comprises two MLP layers, each employing LeakyReLU as the activation function. The parameters of all baseline models are configured according to the default settings specified in the corresponding papers, ensuring consistency in evaluation settings (Ying et al., 2024).
@@ 4.2.2. Ablation Study 1: Shared Critic
Figure 6. Robustness Check of HAFT when confronted with different downstream ML models in terms of F1 score and 1-RAE
An overview of our MARL-based feature selection Framework
@@
Figure 7. Case Study of Messidor_features dataset
An overview of our MARL-based feature selection Framework
@@ 4.2.4. Robustness Check
This experiment aims to answer: Does HAFT exhibit robustness when confronted with various machine learning models serving as downstream tasks? To answer this question, we replace the downstream ML model with Random Forest (RF) (Breiman, 2001), XGBoost (XGB) (Chen and Guestrin, 2016), Support Vector Machine (SVM) (Cortes and Vapnik, 1995), K-Nearest Neighbors (KNN) (Cover and Hart, 1967), and Ridge Regression (Ridge) (Hoerl and Kennard, 1970), respectively. Figure 6 presents the comparison results in terms of F1 score or 1-RAE on Openml_586, Openml_618, Ionosphere, and SpectF datasets. We find that regardless of changes in downstream models, HAFT consistently refines the feature space and achieves better results than the original feature space. The underlying reason is that HAFT can customize feature transformation policies to the specific characteristics of downstream models, resulting in more effective feature spaces. Thus, this experiment demonstrates the robustness of HAFT across different downstream models.
@@ 4.2.8. Study of the Hyperparameter Sensitivity of HAFT
Figure 8. The hyperparameter sensitivity test on SpectF
: Parameter sensitivity
@@ 5. Related Works
Automated Feature Transformation aims to refine or augment the original feature space so that machine learning models can more effectively capture complex, high-order relationships among variables. Traditionally, feature transformation techniques fall into two broad categories: latent representation learning-based and feature transformation-based approaches. Latent representation learning-based methods, as demonstrated in  (Bengio et al., 2013; Guo et al., 2017), focus on learning an optimized, often high-dimensional latent space where feature interactions are implicitly encoded. While these methods are known to boost predictive accuracy, the resulting representations can be challenging to interpret or directly trace back to the original input space. In contrast, feature transformation-based methods (Wang et al., 2022; Huang et al., 2024; Hu et al., 2024; Xiao et al., 2024; Azim et al., 2024) generate new features using arithmetic or aggregate operations. Although these methods yield traceable features, they often encounter an exponentially growing search space problem and become inefficient when dealing with large feature sets. In contrast to previous approaches, our attention-based feature crossing strategy not only captures the inherent relationships among features but also scales efficiently. This is achieved by employing cascading agents guided by a shared critic to learn effective feature interaction policies.
@@
Reinforcement Learning investigates how intelligent agents should act within an environment to maximize the expected cumulative reward (Sutton, 2018; Wiering and Van Otterlo, 2012; Wang et al., 2026). Based on the learned policy, reinforcement learning algorithms can be classified into two main categories: value-based and policy-based. Value-based algorithms (Mnih, 2013; Van Hasselt et al., 2016) estimate the value of states or state-action pairs to facilitate action selection, whereas policy-based algorithms (Sutton et al., 1999) learn a probability distribution that maps states to actions. Furthermore, the actor-critic framework (Schulman et al., 2017) combines the strengths of both approaches. While single-agent reinforcement learning has seen remarkable success, recent work increasingly focuses on multi-agent settings (Wen et al., 2022) where multiple agents collaborate toward a shared goal. However, variations in observations, actions, or rewards often arise, making effective communication critical for coordination and overall performance. Foerster et al. (Foerster et al., 2016) introduced RIAL and DIAL to facilitate communication among agents. Building on this, Foerster et al. (Foerster et al., 2018) proposed the COMA, employing a centralized critic and a counterfactual baseline for credit assignment. This design supports effective coordination even when agents have distinct reward functions. Lowe et al. (Lowe et al., 2017) extended the actor-critic paradigm by implementing a centralized critic for each agent and allowing them to learn continuous policies, improving the scalability and flexibility of MARL systems. Moreover, Lyu et al. (Lyu et al., 2021) found that centralized critics do not consistently outperform their decentralized counterparts across various evaluation domains. Recent advances in the MARL field, such as the theoretical formulation and empirical adaptation of PPO to cooperative multi-agent settings (Yu et al., 2022; Kuba et al., 2021), and Asynchronous Action Coordination (Zhang et al., 2024), have accelerated research progress in this area. Overall, these studies underscore the importance of communication, collaboration, and effective credit assignment in the context of MARL. In our work, we adopt a similar communication mechanism by introducing a customized cascading design alongside a shared critic to address our feature transformation task.